# TAG-WM: Tamper-Aware Generative Image Watermarking via Diffusion Inversion Sensitivity

### v3

@@ 3.3 Localization Performance
Baseline Methods. Previous research [32] has shown that passive methods fail to generalize to unseen tampering types that differ from those encountered during training. For state-of-the-art (SOTA) passive tampering localization methods, such as MVSS-Net [7], OSN [24], PSCC-Net [17], and HiFi-Net [10], the Intersection over Union (IoU) is consistently below 0.3 when tested on unseen optimal image inpainting methods, including ControlNet [30], Stable Diffusion Inpainting [21], and SDXL [19]. Similarly, the Dice score remains below 0.65. Therefore, as a proactive method, we compare our approach exclusively with the SOTA proactive tampering localization method, EditGuard [32], where all tampering types remain unseen during training, with tampering ratios ranging from 0.3 to 0.7.
@@ 3.4 Training Details for Trainable DVRD
Firstly, we constructed the training dataset using 5,000 diverse prompts from the Stable-Diffusion-Prompt dataset. Each prompt was used to generate an image with the Stable Diffusion v2.1 model under five different ODE schedulers—DDIM [22], UniPC [34], PNDM [16], DEIS [31], and DPMSolver [18]—resulting in a total of 5,000 images (1,000 images per scheduler). Secondly, Each generated image underwent all of the following tampering types: random cropping (crop ratio=0.1, 0.3, 0.5, 0.7 and 0.9), random pixel dropping (drop ratio=0.1, 0.3, 0.5, 0.7 and 0.9), random logo insertion with varying numbers and sizes (logo count-logo ratio=1-0.7, 3-0.39, 5-0.25, 7-0.2 and 9-0.1), and image inpainting using MAT [15] pre-trained on three datasets—CelebA-HQ [^2] https://www.kaggle.com/datasets/vincenttamml/celebamaskhq512, FFHQ [^3] https://huggingface.co/datasets/LIAGM/FFHQ_datasets, and Places365-Standard [^4] https://paperswithcode.com/dataset/places365—all at a resolution of $512\times 512$. These tampering techniques simulate a variety of altered regions with different shapes, sizes, and edge smoothness. Note that our proactive tampering localization method does not require distinguishing between tampering types; thus, the aforementioned manipulations are sufficient for our purposes. For each tampered image, we also stored the corresponding ground truth tampering mask. Finally, we split the dataset into training and validation sets at a ratio of $0.95:0.05$.

### v4

@@ 5.3 Localization Performance
Baseline Methods. Previous research [wm_and_tl-Editguard] has shown that passive methods fail to generalize to unseen tampering types that differ from those encountered during training. For state-of-the-art (SOTA) passive tampering localization methods, such as MVSS-Net [tl-passive-universal-MVSSNet], OSN [tl-passive-specific-OSN], PSCC-Net [tl-passive-universal-PSCCNet], and HiFi-Net [tl-passive-universal-HiFiNet], the Intersection over Union (IoU) is consistently below 0.3 when tested on unseen optimal image inpainting methods, including ControlNet [ImageInpainting-ControlNet], Stable Diffusion Inpainting [DMs-LDMs], and SDXL [ImageInpainting-SDXL]. Similarly, the Dice score remains below 0.65. Therefore, as a proactive method, we compare our approach exclusively with the SOTA proactive tampering localization method, EditGuard [wm_and_tl-Editguard], where all tampering types remain unseen during training, with tampering ratios ranging from 0.3 to 0.7.
@@ 5.4 Training Details for Trainable DVRD
Firstly, we constructed the training dataset using 5,000 diverse prompts from the Stable-Diffusion-Prompt dataset. Each prompt was used to generate an image with the Stable Diffusion v2.1 model under five different ODE schedulers—DDIM [DMs-scheduler-DDIM], UniPC [DMs-scheduler-UniPC], PNDM [DMs-scheduler-PNDM], DEIS [DMs-scheduler-DEIS], and DPMSolver [DMs-scheduler-DPMSolver]—resulting in a total of 5,000 images (1,000 images per scheduler). Secondly, Each generated image underwent all of the following tampering types: random cropping (crop ratio=0.1, 0.3, 0.5, 0.7 and 0.9), random pixel dropping (drop ratio=0.1, 0.3, 0.5, 0.7 and 0.9), random logo insertion with varying numbers and sizes (logo count-logo ratio=1-0.7, 3-0.39, 5-0.25, 7-0.2 and 9-0.1), and image inpainting using MAT [ImageInpainting-MAT] pre-trained on three datasets—CelebA-HQ [^2] https://www.kaggle.com/datasets/vincenttamml/celebamaskhq512, FFHQ [^3] https://huggingface.co/datasets/LIAGM/FFHQ_datasets, and Places365-Standard [^4] https://paperswithcode.com/dataset/places365—all at a resolution of $512\times 512$. These tampering techniques simulate a variety of altered regions with different shapes, sizes, and edge smoothness. Note that our proactive tampering localization method does not require distinguishing between tampering types; thus, the aforementioned manipulations are sufficient for our purposes. For each tampered image, we also stored the corresponding ground truth tampering mask. Finally, we split the dataset into training and validation sets at a ratio of $0.95:0.05$.